# QMill: Representative Quantum Data Generation for Quantum Machine Learning Utility

### v2

@@ Abstract
Quantum machine learning (QML) promises significant speedups, particularly when operating on quantum datasets. However, its progress is hindered by the scarcity of suitable training data. Existing synthetic data generation methods fall short in capturing essential entanglement properties, limiting their utility for QML. To address this, we introduce QMill, a low-depth quantum data generation framework that produces entangled, high-quality samples emulating diverse classical and quantum distributions, enabling more effective development and evaluation of QML models in representative-data settings.
@@ 1 Introduction
Quantum machine learning (QML) is emerging as a transformative field, with applications ranging from image recognition to scientific computing (Riofrio et al., 2024; Liang et al., 2023; Peral-García et al., 2024; Wang et al., 2022; Guan et al., 2021). QML offers theoretical speedups over classical methods—but crucially, these speedups are provably attainable when operating on quantum datasets, i.e., data exhibiting superposition, interference, and entanglement (Biamonte et al., 2017; Carleo et al., 2019; DiBrita et al., 2024; Beaudoin et al., 2022; Hu et al., 2022; Delgado & Hamilton, 2022). Despite this, nearly all existing QML research focuses on classical data inputs due to the scarcity of real-world quantum datasets (Silver et al., 2022; 2023b). Quantum data is difficult to obtain: current quantum sensing technology is nascent, measurements are inherently probabilistic, large-scale data collection is cost-prohibitive, and noise from environmental and control sources further limits usability (Degen et al., 2017; Aslam et al., 2023). This data gap has become a fundamental bottleneck preventing the community from developing and validating QML models that can operate directly on quantum data, the very regime where QML promises a provable advantage.
@@
To address these challenges, we present QMill, a versatile quantum data generation framework designed to produce synthetic datasets that reflect diverse CE distributions and faithfully emulate both classical and quantum structures.
@@
  • QMill generates synthetic datasets that capture a range of concentratable entanglement (CE) values, reflecting the variability observed in real-world quantum data.
  • We design low-depth ansatzes tailored to Gaussian, Weibull, and Uniform distributions, enabling QMill to stress-test statistical behavior under quantum constraints.
  • By leveraging dual annealing (Sahin & Ciric, 1998), QMill optimizes entangled states efficiently, ensuring compatibility with contemporary quantum hardware.
  • QMill incorporates SWAP tests (Zhang et al., 2024) to guarantee sample diversity and reduce redundancy, crucial for training generalizable QML models.
  • We demonstrate QMill’s versatility across classical datasets (e.g., MNIST (Deng, 2012), FashionMNIST (Xiao et al., 2017), CIFAR-10 (Krizhevsky et al., 2009)) and quantum datasets (e.g., quantum chemistry (Perrier et al., 2022), soil moisture (Arumugam et al., 2024), dark matter (Chen et al., 2024)), achieving a deviation of <0.1<0.1 from the target CE distributions.
  • To show QMill’s practical utility, we train a quantum neural network on QMill-generated CE feature sets and show an 84.8% accuracy against a classical ceiling.
  • QMill’s data generation methodology, machine learning codebase, and generated datasets are open-sourced at: https://github.com/positivetechnologylab/QMill.
@@ 2.2 Variational Quantum Circuits and Noise
Variational quantum circuits (VQCs), or ansatzes, are widely used in QML due to their tunability and expressiveness (Wang et al., 2022; DiBrita et al., 2024; Han et al., 2025). Each gate in a VQC is parameterized (e.g., $R_{y}(\theta)$), and the overall state $|\psi(\vec{\theta})\rangle=U(\vec{\theta})|\psi_{0}\rangle$ depends on a set of parameters $\vec{\theta}$. These parameters are optimized to minimize a classical loss function $f(\vec{\theta})$. On real hardware, especially NISQ devices, gate noise is a key challenge. As each gate has a non-zero error rate $\epsilon$, the total error grows with depth $d$ approximately as $1-(1-\epsilon)^{d}$ (Silver et al., 2023a; Bhattacharjee et al., 2019; Ash-Saki et al., 2019). Shallow circuits are therefore crucial to maintain high fidelity, especially for QML tasks. QMill leverages low-depth ansatzes to mitigate this noise while preserving expressivity.
@@ 2.4 Concentratable Entanglement (CE)
3 Motivation for QMill
Progress in QML is held back by the lack of scalable, diverse, and entanglement-aware quantum datasets. Existing quantum datasets are small and expensive to generate, and current synthetic methods are even more limited (Zoufal et al., 2019; Benedetti et al., 2019). The most notable effort by Schatzki et al. (Schatzki et al., 2021) proposes training ansatzes to match a fixed CE target, but their approach often fails to reach the desired CE value and ignores a more fundamental issue: real quantum data does not have a single entanglement level. In practice, quantum datasets exhibit a spread of CE values across samples. Training and benchmarking QML models on a fixed CE setting oversimplifies the problem and leads to poor generalization. What is needed instead is a generator that can produce datasets with controlled CE distributions – capturing the full range from weak to strong entanglement. QMill fills this gap. It generates synthetic datasets where CE values follow user-specified distributions. It uses low-depth, distribution-specific ansatzes optimized via annealing methods, making it both noise-resilient and efficient. The result is a scalable framework for producing entanglement-rich, diverse, and realistic quantum datasets, enabling the next stage of data-driven QML development.
4 QMill’s Design
Figure 1: QMill takes classical product states and generates diverse and customizable quantum data for QML tasks.
QMill is a quantum data generation framework designed to create high-quality synthetic datasets for QML. Its core goal is to generate entangled states that match a target distribution of concentratable entanglement (CE) while remaining shallow enough to run on noisy hardware. As shown in Fig. 1, QMill starts with Haar-random product states, applies a parameterized ansatz to entangle them, optimizes the circuit to match a CE distribution, and validates sample diversity in the generated dataset via SWAP tests.
The framework has four components: (A) a set of low-depth variational ansatzes supporting different entanglement structures, (B) a pipeline for sample generation and CE measurement using efficient density matrix approximations, (C) a dual-annealing optimization loop minimizing CE distribution mismatch, and (D) a SWAP test–based diversity check to avoid mode collapse. Together, these components make QMill scalable, customizable, and hardware-compatible.
@@ 4.1 Parameterized Circuits & Objective Function
Figure 2: QMill develops a variety of ansatz designs for real and synthetic CE distributions.
The main design tension is between expressibility and hardware feasibility: deeper circuits can model richer CE distributions, but are more prone to noise on near-term hardware. To explore this trade-off, QMill includes four low-depth parameterized circuits (A1–A4), shown in Fig. 2, each probing different entanglement and noise behaviors. A1 uses compact RX, RZ, and controlled-RZ gates. A2 extends A1 with a denser entangling pattern. A3 incorporates Hadamard and controlled-RX gates. A4 combines RX, RY, CNOT, and controlled-RZ gates in a deeper structure.
The goal is not to find a universal best ansatz, but to evaluate which structures best match the target CE under depth constraints. Parameters $\vec{\theta}$ are tuned using dual annealing (Sahin & Ciric, 1998), a global optimizer effective in non-convex landscapes where gradient methods often fail, especially for skewed or multimodal CE targets. The objective is to minimize the total variation distance (TVD) between the empirical CE histogram and the target:
@@ 4.2 Sample Generation and CE Measurement
To measure CE, QMill considers an efficient approximation from Beckey et al. (Beckey et al., 2021):
@@ 4.3 Concentratable Entanglement Distributions
Figure 3: In addition to the CE distributions of real data, QMill also tests its efficacy for different CE distributions.
A core strength of QMill is its ability to match full distributions of CE values, not just a single entanglement target. This is essential because real quantum datasets rarely have uniform entanglement; instead, they exhibit broad or skewed CE profiles. Supporting full CE distributions enables realistic benchmarking of QML models across diverse entanglement regimes. QMill supports both real and synthetic targets. For real CE distributions, we extract histograms from quantum-encoded classical datasets such as MNIST, FashionMNIST, and CIFAR-10 (Krizhevsky et al., 2009; Xiao et al., 2017; Deng, 2012), as well as native quantum datasets like quantum chemistry, soil moisture, and dark matter (Arumugam et al., 2024; Chen et al., 2024; Schütt et al., 2017). Each dataset is amplitude encoded, and CE is computed to produce empirical histograms used as generation targets. To stress-test QMill’s flexibility, we define several synthetic CE distributions:
@@
Fig. 3 shows target examples. During training, QMill bins CE values from generated samples and compares them to the target via TVD. This approach allows controlled exploration of how QML models respond to various entanglement regimes. For instance, one can test how ansatz performance varies under low vs. high CE, or compare classical and quantum dataset demands. QMill thus enables entanglement-aware dataset engineering, which comprises more than just data generation.
Figure 4: QMill uses the SWAP test to validate the dissimilarity of any two random samples with similar CE values.
@@ 4.4 SWAP Test for Sample Diversity Validation
Matching CE distributions alone doesn’t guarantee dataset quality. A generator could produce near-identical states with the same CE, resulting in low diversity and poor generalization. Ensuring that QMill outputs not just entangled, but distinct samples is therefore critical. To enforce diversity, QMill uses the SWAP test (Zhang et al., 2024) (see App. A.2 for details), a quantum routine that measures the fidelity between two states:
@@
High fidelity ($\approx 1$) indicates similarity; values near $0.5$ suggest dissimilarity. Unlike classical similarity checks, the SWAP test is efficient and non-destructive. During training, a random subset of sample pairs is selected, and their average SWAP test score is calculated. If average fidelity exceeds a threshold (e.g., >0.95>0.95), this signals mode collapse. In response, QMill introduces a diversity penalty to steer optimization away from redundant states, especially important for sharp or skewed CE targets. By combining CE alignment with active diversity enforcement, QMill produces datasets that are both representative of the target entanglement structure and richly varied at the state level to generate diverse dataset samples.
In summary, QMill unifies low-depth ansatzes, CE-targeted optimization, and diversity validation into a practical pipeline for generating high-quality quantum datasets. Each component addresses a key challenge—circuit noise, CE fidelity, and sample uniqueness—resulting in a scalable framework ready for QML training and evaluation.
5 QMill’s Implementation and Methods
@@ 5.1 Experimental and Software Setup
We evaluate QMill using Qiskit Aer’s circuit simulator with IBM Sherbrooke’s noise model for noisy simulations. Real-machine experiments are also conducted on IBM Sherbrooke. All circuits are implemented in Python 3.10.12 using Qiskit 1.2 (Aleksandrowicz et al., 2019). Simulations are executed on a local research cluster running Ubuntu 22.04.2 LTS, with a 32-core 2.0 GHz AMD EPYC 7551P processor and 32 GB RAM. Each experiment uses 2048 measurement shots. Circuits are constructed using Qiskit’s QuantumCircuit class, and noiseless simulations are performed for baseline evaluations.
@@ 5.2 Evaluated Classical and Quantum Datasets
To evaluate QMill’s ability to generate quantum data with controlled CE characteristics, we use both synthetic and real datasets. For stress-testing, we define four synthetic CE target distributions over the interval [0, 0.4]. These include a uniform distribution, a Gaussian distribution centered at 0.2 with a standard deviation of 0.05, a left-skewed Weibull distribution (shape parameter 1.2, scaled by 0.05), and a right-skewed variant obtained by reflecting the left-skewed version across $x=0.2$. These distributions are chosen to span a wide range of entanglement behaviors observed in real quantum systems. We also evaluate CE profiles derived from classical datasets: MNIST (Deng, 2012), FashionMNIST (Xiao et al., 2017), and CIFAR-10 (Krizhevsky et al., 2009).
The data are standardized and reduced in dimension using Principal Component Analysis (PCA) (Pearson, 1901) to $2^{n}-1$ features for $n$ qubits. These features are then embedded into quantum amplitudes using amplitude encoding (Rath & Date, 2024). The resulting quantum states are processed to compute CE values as described in Sec. 4, and their empirical CE distributions are scaled for comparison against QMill outputs. More significantly, we evaluate CE targets extracted from three quantum datasets. The quantum chemistry dataset (Perrier et al., 2022) contains 134k molecules from QM9, each represented using engineered features derived from atomic and molecular statistics. These include atomic charge moments, vibrational frequencies, spatial metrics, and element counts, all aggregated into fixed-length vectors suitable for amplitude encoding.
For quantum-sensed workloads, we simulate two protocols. The first is a soil moisture sensing setup based on the STQS framework (Jebraeilli et al., 2025; Arumugam et al., 2024), which uses entangled Rydberg atoms to sense phase differences from soil reflections. Simulations are run for both high and low moisture regimes, incorporating phase jitter to generate ensembles of quantum states. CE values are computed for each state to form CE distributions reflective of different sensing environments. The second protocol is a dark matter detection setup adapted from (Chen et al., 2024), using a four-qubit sensing circuit where the signal strength $\phi$ encodes the dark matter interaction. Simulations with phi=0.01\ phi=0.01 and phi=0.1\ phi=0.1 yield distinct CE distributions, enabling us to evaluate QMill under both weak and strong signal conditions. See App. B for sensor circuit details.
5.3 QMill’s Evaluation Metrics

### v3

# Generating Data with Quantum Properties for Quantum Machine Learning Utility

@@ Abstract
Quantum machine learning (QML) promises significant speedups, particularly when operating on quantum datasets. However, its progress is hindered by the scarcity of suitable training data. Existing synthetic data generation methods fall short in capturing essential entanglement properties, limiting their utility for QML. To address this, we introduce QuFoundry, a low-depth quantum data generation framework that produces entangled, high-quality samples emulating diverse classical and quantum distributions, enabling more effective development and evaluation of QML models in representative-data settings.
@@ 1 Introduction
Quantum machine learning (QML) is emerging as a transformative field, with applications ranging from image recognition to scientific computing (Riofrio et al., 2024; Liang et al., 2023; Peral-García et al., 2024; Wang et al., 2022; Guan et al., 2021). QML offers theoretical speedups over classical methods—but crucially, these speedups are provably attainable when operating on quantum datasets, i.e., data exhibiting superposition, interference, and entanglement (Biamonte et al., 2017; Carleo et al., 2019; DiBrita et al., 2024; Beaudoin et al., 2022; Hu et al., 2022; Delgado and Hamilton, 2022). Despite this, nearly all existing QML research focuses on classical data inputs due to the scarcity of real-world quantum datasets (Silver et al., 2022; 2023b). Quantum data is difficult to obtain: current quantum sensing technology is nascent, measurements are inherently probabilistic, large-scale data collection is cost-prohibitive, and noise from environmental and control sources further limits usability (Degen et al., 2017; Aslam et al., 2023). This data gap has become a fundamental bottleneck preventing the community from developing and validating QML models that can operate directly on quantum data, the very regime where QML promises a provable advantage.
@@
To address these challenges, we present QuFoundry, a versatile quantum data generation framework designed to produce synthetic datasets that reflect diverse CE distributions and faithfully emulate both classical and quantum structures.
@@
  • QuFoundry generates synthetic datasets that capture a range of concentratable entanglement (CE) values, reflecting the variability observed in real-world quantum data.
  • We design low-depth ansatzes tailored to Gaussian, Weibull, and Uniform distributions, enabling QuFoundry to stress-test statistical behavior under quantum constraints.
  • By leveraging dual annealing (Sahin and Ciric, 1998), QuFoundry optimizes entangled states efficiently, ensuring compatibility with contemporary quantum hardware.
  • QuFoundry incorporates SWAP tests (Zhang et al., 2024) to guarantee sample diversity and reduce redundancy, crucial for training generalizable QML models.
  • We demonstrate QuFoundry’s versatility across classical datasets (e.g., MNIST (Deng, 2012), FashionMNIST (Xiao et al., 2017), CIFAR-10 (Krizhevsky et al., 2009)) and quantum datasets (e.g., quantum chemistry (Perrier et al., 2022), soil moisture (Arumugam et al., 2024), dark matter (Chen et al., 2024)), achieving a deviation of <0.1<0.1 from the target CE distributions.
  • To show QuFoundry’s practical utility, we train a quantum neural network on QuFoundry-generated CE feature sets and show an 84.8% accuracy against a classical ceiling.
  • QuFoundry’s data generation methodology, machine learning codebase, and generated datasets are open-sourced at: https://github.com/positivetechnologylab/QuFoundry.
@@ 2.2 Variational Quantum Circuits and Noise
Variational quantum circuits (VQCs), or ansatzes, are widely used in QML due to their tunability and expressiveness (Wang et al., 2022; DiBrita et al., 2024; Han et al., 2025). Each gate in a VQC is parameterized (e.g., $R_{y}(\theta)$), and the overall state $|\psi(\vec{\theta})\rangle=U(\vec{\theta})|\psi_{0}\rangle$ depends on a set of parameters $\vec{\theta}$. These parameters are optimized to minimize a classical loss function $f(\vec{\theta})$. On real hardware, especially NISQ devices, gate noise is a key challenge. As each gate has a non-zero error rate $\epsilon$, the total error grows with depth $d$ approximately as $1-(1-\epsilon)^{d}$ (Silver et al., 2023a; Bhattacharjee et al., 2019; Ash-Saki et al., 2019). Shallow circuits are therefore crucial to maintain high fidelity, especially for QML tasks. QuFoundry leverages low-depth ansatzes to mitigate this noise while preserving expressivity.
@@ 2.4 Concentratable Entanglement (CE)
3 Motivation for QuFoundry
Progress in QML is held back by the lack of scalable, diverse, and entanglement-aware quantum datasets. Existing quantum datasets are small and expensive to generate, and current synthetic methods are even more limited (Zoufal et al., 2019; Benedetti et al., 2019). The most notable effort by Schatzki et al. (Schatzki et al., 2021) proposes training ansatzes to match a fixed CE target, but their approach often fails to reach the desired CE value and ignores a more fundamental issue: real quantum data does not have a single entanglement level. In practice, quantum datasets exhibit a spread of CE values across samples. Training and benchmarking QML models on a fixed CE setting oversimplifies the problem and leads to poor generalization. What is needed instead is a generator that can produce datasets with controlled CE distributions – capturing the full range from weak to strong entanglement. QuFoundry fills this gap. It generates synthetic datasets where CE values follow user-specified distributions. It uses low-depth, distribution-specific ansatzes optimized via annealing methods, making it both noise-resilient and efficient. The result is a scalable framework for producing entanglement-rich, diverse, and realistic quantum datasets, enabling the next stage of data-driven QML development.
4 QuFoundry’s Design
Figure 1: QuFoundry takes classical product states and generates diverse and customizable quantum data for QML tasks.
QuFoundry is a quantum data generation framework designed to create high-quality synthetic datasets for QML. Its core goal is to generate entangled states that match a target distribution of concentratable entanglement (CE) while remaining shallow enough to run on noisy hardware. As shown in Fig. 1, QuFoundry starts with Haar-random product states, applies a parameterized ansatz to entangle them, optimizes the circuit to match a CE distribution, and validates sample diversity in the generated dataset via SWAP tests.
The framework has four components: (A) a set of low-depth variational ansatzes supporting different entanglement structures, (B) a pipeline for sample generation and CE measurement using efficient density matrix approximations, (C) a dual-annealing optimization loop minimizing CE distribution mismatch, and (D) a SWAP test–based diversity check to avoid mode collapse. Together, these components make QuFoundry scalable, customizable, and hardware-compatible.
@@ 4.1 Parameterized Circuits & Objective Function
Figure 2: QuFoundry develops a variety of ansatz designs for real and synthetic CE distributions.
The main design tension is between expressibility and hardware feasibility: deeper circuits can model richer CE distributions, but are more prone to noise on near-term hardware. To explore this trade-off, QuFoundry includes four low-depth parameterized circuits (A1–A4), shown in Fig. 2, each probing different entanglement and noise behaviors. A1 uses compact RX, RZ, and controlled-RZ gates. A2 extends A1 with a denser entangling pattern. A3 incorporates Hadamard and controlled-RX gates. A4 combines RX, RY, CNOT, and controlled-RZ gates in a deeper structure.
The goal is not to find a universal best ansatz, but to evaluate which structures best match the target CE under depth constraints. Parameters $\vec{\theta}$ are tuned using dual annealing (Sahin and Ciric, 1998), a global optimizer effective in non-convex landscapes where gradient methods often fail, especially for skewed or multimodal CE targets. The objective is to minimize the total variation distance (TVD) between the empirical CE histogram and the target:
@@ 4.2 Sample Generation and CE Measurement
To measure CE, QuFoundry considers an efficient approximation from Beckey et al. (Beckey et al., 2021):
@@ 4.3 Concentratable Entanglement Distributions
Figure 3: In addition to the CE distributions of real data, QuFoundry also tests its efficacy for different CE distributions.
A core strength of QuFoundry is its ability to match full distributions of CE values, not just a single entanglement target. This is essential because real quantum datasets rarely have uniform entanglement; instead, they exhibit broad or skewed CE profiles. Supporting full CE distributions enables realistic benchmarking of QML models across diverse entanglement regimes. QuFoundry supports both real and synthetic targets. For real CE distributions, we extract histograms from quantum-encoded classical datasets such as MNIST, FashionMNIST, and CIFAR-10 (Krizhevsky et al., 2009; Xiao et al., 2017; Deng, 2012), as well as native quantum datasets like quantum chemistry, soil moisture, and dark matter (Arumugam et al., 2024; Chen et al., 2024; Schütt et al., 2017). Each dataset is amplitude encoded, and CE is computed to produce empirical histograms used as generation targets. To stress-test QuFoundry’s flexibility, we define several synthetic CE distributions:
@@
Fig. 3 shows target examples. During training, QuFoundry bins CE values from generated samples and compares them to the target via TVD. This approach allows controlled exploration of how QML models respond to various entanglement regimes. For instance, one can test how ansatz performance varies under low vs. high CE, or compare classical and quantum dataset demands. QuFoundry thus enables entanglement-aware dataset engineering, which comprises more than just data generation.
Figure 4: QuFoundry uses the SWAP test to validate the dissimilarity of any two random samples with similar CE values.
@@ 4.4 SWAP Test for Sample Diversity Validation
Matching CE distributions alone doesn’t guarantee dataset quality. A generator could produce near-identical states with the same CE, resulting in low diversity and poor generalization. Ensuring that QuFoundry outputs not just entangled, but distinct samples is therefore critical. To enforce diversity, QuFoundry uses the SWAP test (Zhang et al., 2024) (see App. A.2 for details), a quantum routine that measures the fidelity between two states:
@@
High fidelity ($\approx 1$) indicates similarity; values near $0.5$ suggest dissimilarity. Unlike classical similarity checks, the SWAP test is efficient and non-destructive. During training, a random subset of sample pairs is selected, and their average SWAP test score is calculated. If average fidelity exceeds a threshold (e.g., >0.95>0.95), this signals mode collapse. In response, QuFoundry introduces a diversity penalty to steer optimization away from redundant states, especially important for sharp or skewed CE targets. By combining CE alignment with active diversity enforcement, QuFoundry produces datasets that are both representative of the target entanglement structure and richly varied at the state level to generate diverse dataset samples.
In summary, QuFoundry unifies low-depth ansatzes, CE-targeted optimization, and diversity validation into a practical pipeline for generating high-quality quantum datasets. Each component addresses a key challenge—circuit noise, CE fidelity, and sample uniqueness—resulting in a scalable framework ready for QML training and evaluation.
5 QuFoundry’s Implementation and Methods
@@ 5.1 Experimental and Software Setup
We evaluate QuFoundry using Qiskit Aer’s circuit simulator with IBM Sherbrooke’s noise model for noisy simulations. Real-machine experiments are also conducted on IBM Sherbrooke. All circuits are implemented in Python 3.10.12 using Qiskit 1.2 (Aleksandrowicz et al., 2019). Simulations are executed on a local research cluster running Ubuntu 22.04.2 LTS, with a 32-core 2.0 GHz AMD EPYC 7551P processor and 32 GB RAM. Each experiment uses 2048 measurement shots. Circuits are constructed using Qiskit’s QuantumCircuit class, and noiseless simulations are performed for baseline evaluations.
@@ 5.2 Evaluated Classical and Quantum Datasets
To evaluate QuFoundry’s ability to generate quantum data with controlled CE characteristics, we use both synthetic and real datasets. For stress-testing, we define four synthetic CE target distributions over the interval [0, 0.4]. These include a uniform distribution, a Gaussian distribution centered at 0.2 with a standard deviation of 0.05, a left-skewed Weibull distribution (shape parameter 1.2, scaled by 0.05), and a right-skewed variant obtained by reflecting the left-skewed version across $x=0.2$. These distributions are chosen to span a wide range of entanglement behaviors observed in real quantum systems. We also evaluate CE profiles derived from classical datasets: MNIST (Deng, 2012), FashionMNIST (Xiao et al., 2017), and CIFAR-10 (Krizhevsky et al., 2009).
The data are standardized and reduced in dimension using Principal Component Analysis (PCA) (Pearson, 1901) to $2^{n}-1$ features for $n$ qubits. These features are then embedded into quantum amplitudes using amplitude encoding (Rath and Date, 2024). The resulting quantum states are processed to compute CE values as described in Sec. 4, and their empirical CE distributions are scaled for comparison against QuFoundry outputs. More significantly, we evaluate CE targets extracted from three quantum datasets. The quantum chemistry dataset (Perrier et al., 2022) contains 134k molecules from QM9, each represented using engineered features derived from atomic and molecular statistics. These include atomic charge moments, vibrational frequencies, spatial metrics, and element counts, all aggregated into fixed-length vectors suitable for amplitude encoding.
For quantum-sensed workloads, we simulate two protocols. The first is a soil moisture sensing setup based on the STQS framework (Jebraeilli et al., 2025; Arumugam et al., 2024), which uses entangled Rydberg atoms to sense phase differences from soil reflections. Simulations are run for both high and low moisture regimes, incorporating phase jitter to generate ensembles of quantum states. CE values are computed for each state to form CE distributions reflective of different sensing environments. The second protocol is a dark matter detection setup adapted from (Chen et al., 2024), using a four-qubit sensing circuit where the signal strength $\phi$ encodes the dark matter interaction. Simulations with phi=0.01\ phi=0.01 and phi=0.1\ phi=0.1 yield distinct CE distributions, enabling us to evaluate QuFoundry under both weak and strong signal conditions. See App. B for sensor circuit details.
5.3 QuFoundry’s Evaluation Metrics